# RS-ISRefiner: Towards Better Adapting Vision Foundation Models for Interactive Segmentation of Remote Sensing Images

### v1

@@ III-C High-Quality Feature Extractor
While the frozen VFM backbone provides rich global semantics and structural priors, these features alone are insufficient to capture fine-grained local variations in remote sensing targets. Therefore, we design a High-Quality Feature Extractor to complement the VFM outputs with detailed spatial representations. This module consists of two components, namely the Image Feature Exactor and the Image Feature Fusion, as shown in Fig.4.
Figure 4: High Quality Feature Extractor. This module comprises an image feature extractor (light-colored region in the middle) and a feature fusion module (green region on the right). The image feature extractor includes one ResBlock, three DeformableResBlocks, and one 1×1 convolution, with the detailed architectures of ResBlock and Deformable ResBlock shown on the left of the diagram. The feature fusion module is primarily composed of cross-attention blocks, where $\theta$ serves as a learnable parameter.
The Image Feature Extractor adopts a residual network architecture, comprising one convolutional residual module, three deformable convolutional residual modules and one linear projection layer. Herein, deformable convolutions are introduced to learn sampling position offsets, enabling adaptive adaptation to variations in the shape, orientation and scale of remote sensing targets. Moreover, the integration of the Squeeze-and-Excitation Attention mechanism into the deformable convolutional residual modules servers as a commonly employed strategy in convolutional networks to enhance the feature extraction capability of the network. Upon processing image $I$ through the Image Feature Extractor, a spatially informative feature $F_{IFE}$ is derived.
@@
Finally, the fused feature $F_{IFE}^{\prime}$ undergoes shape adjustment and is unsampled to a high resolution, yielding the high-quality image feature $F_{HQ}$.
@@ III-D Transformer Decoder
Built upon the high-level semantic features extracted by the frozen Vision Foundation Model, the decoder establishes effective information interaction between the VFM’s global representations and the locally enhanced features from the extractor module. Through bidirectional cross-attention processing, the model establishes efficient information interaction between global semantics and local details, thereby generating prediction results with both global consistency and high-precision boundaries, as illustrated in Fig.5.
First, $F_{mask}$ and $F_{ViT}^{N}$ are concatenated, then undergo channel-wise fusion via the Dense Fusion Convolution, which primarily consists of $1\times 1$ convolution operations, to align dimensions with those of token $Q$. The resulting features $F_{DFC}$, in conjunction with $Q$, are fed into a two-layer bidirectional cross-attention module. Diverging from the conventional Transformer Decoder architecture, we herein reverse the sequence of token self-attention and cross-attention operations. As highlighted in [50], tokens are initially randomly initialized and devoid of image feature information, rendering self-attention operations ineffective for acquiring sufficient information at the initial stage. Thus, the token-to-feature cross-attention operation is performed. $Q$ serves as the Query, with Keys and Values derived from $F_{DFC}$, to align $Q$ with image features. The output then undergoes residual connection and normalization processing. The formula is as follows.
@@ III-E Probability Map Modulation
As noted in [18], probability map modulation enhances or suppresses the probability values of local regions based on the probability map obtained from the previous round and the current user’s click information, making them closer to the foreground or background distribution desired by the user. However, in the original paper, the modulated local region is a circular area, where both foreground and background pixels may coexist. During the probability modification process, a positive click is intended to enhance the foreground probability, but since there is no distinction between foreground and background within this circular region, the background probability is also inadvertently enhanced. To address this issue, we propose an improved probability map modulation algorithm, as illustrated in Fig.6.
Figure 6: Modulate Map Mechanism. The left image shows the uncorrected result, while the right image presents the corrected version. The modualtion window (gray region in the middle image) is extracted using the formula on the left, and the pixel values within this window are refined according to the formula on the right.
After the model outputs the predicted probability map, we obtain the latest click point $u=(x,y,t)\in C$ in the current iteration, where $x$ and $y$ represent the horizontal and vertical coordinates, respectively, and $t$ denotes the click type, with $t=1$ indicating a foreground click and $t=0$ indicating a background click. The modulation radius $R$ is then determined based on the click point: if an opposite-type click exists in the vicinity, $R$ is set to half of the distance from the current click to the nearest opposite-type click; otherwise, a predefined radius $R_{max}$ is used. This design prevents the modulation regions of foreground and background from overlapping. To avoid an excessively small modulation radius, we also introduce a minimum radius constraint $R_{min}$. The equation is given as follows.
@@
where $O$ denotes the set of clicks of opposite type. After obtaining the modulation radius, a circular modulation window is constructed with the click point as the center, which is the approach adopted in the original paper. In this paper, considering that the foreground and background pixel probabilities in the probability map exhibit more or less differences, we introduce conditional filtering to sieve out some pixels and extract only the regions belonging to the foreground or background as the modulation window, instead of modulating the entire circular area. The conditional filtering formula is as follows.
@@
herein, $p_{x\in W}$ denotes the pixel probability within the circular region, $mean(\cdot)$ represents the average value of the probability pixels in the region, $mid(\cdot)$ refers to the median of the probability pixels in the region, and $p_{u}$ denotes the pixel probability at the click point. Depending on the click type $t_{u}$, pixels whose probability values fall between the minimum and maximum screening thresholds are retained. Subsequently, the $\gamma$ modulation follows the approach of the original paper: for different click types, the pixel probabilities in the probability map are adjusted, with the adjustment magnitude decreasing as the distance from the click point increases. The relevant formulas are as follows.
@@
[…]
where $d_{x}=\left\|x-u\right\|$, represents the distance between the pixel $x$ and the click $u$. The modulated probability map will serve as the input for the next iteration of the model.
TABLE I: Comparison of our method with state-of-the-art approaches in NoC80, NoC85 and NoC90 on the iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding datasets. The best results are highlighted in bold, while the second-best are underlined. * denotes that the backbone parameters of SAM in SAM-HQ are frozen. ** indicates that PiClick was trained without multi-output prediction, using the same model and loss function as the original PiClick.
Method Backbone iSAID SandBar NWPU LoveDA Urban WHUBuilding NoC80 NoC85 NoC90 NoC80 NoC85 NoC90 NoC80 NoC85 NoC90 NoC80 NoC85 NoC90 NoC80 NoC85 NoC90 RITM HRNet32 2.37 3.74 6.71 5.11 8.29 13.54 3.25 4.51 7.17 6.73 8.36 10.82 3.82 5.41 9.00 CDNet ResNet34 5.47 8.14 12.09 8.45 11.76 16.06 5.30 7.84 10.57 9.73 12.58 16.17 7.36 10.20 14.54 FocalClick HRNet32 6.14 8.28 12.14 8.78 11.85 15.29 5.74 7.78 10.70 9.24 11.84 15.14 7.51 10.06 14.24 GPCIS ResNet50 3.78 5.84 9.66 8.35 11.34 15.09 3.87 5.67 8.77 7.00 9.18 12.51 5.48 7.65 11.77 SAM-HQ* ViT-B 10.70 13.47 16.71 11.14 14.98 18.29 11.63 13.31 15.31 15.95 17.71 19.15 11.23 14.40 18.13 SimpleClick ViT-B 2.40 3.76 6.66 5.25 8.22 13.34 3.08 4.57 7.12 6.82 8.33 10.62 3.64 5.08 8.42 PiClick** ViT-B 3.01 4.61 8.28 5.42 8.64 14.06 4.12 6.06 8.61 8.31 10.16 12.85 3.87 5.35 8.98 AdaptiveClick ViT-B 2.89 4.61 7.79 6.29 9.37 13.76 3.72 5.36 7.81 6.84 8.44 10.67 3.54 5.03 8.46 MST ViT-B 2.77 4.40 8.20 6.15 9.55 14.19 3.08 4.75 7.71 6.84 8.72 11.60 4.16 6.17 10.55 MFP ViT-B 2.20 3.46 6.62 4.63 7.91 12.97 3.24 4.56 7.30 7.13 8.79 11.62 4.01 5.63 9.28 Ours ViT-B 2.25 3.38 6.51 4.54 7.67 12.90 2.53 3.83 6.56 5.99 7.62 10.19 3.65 5.03 8.42
TABLE II: Comparison of our method against state-of-the-art approaches in NoF80, NoF85 and NoF90 on the iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding datasets. The best results are shown in bold and the second-best are underlined. The symbols * and ** follow the same definitions as in Table I.
Method Backbone iSAID 844 SandBar 663 NWPU 2049 LoveDA Urban 851 WHUBuilding 2559 NoF80 NoF85 NoF90 NoF80 NoF85 NoF90 NoF80 NoF85 NoF90 NoF80 NoF85 NoF90 NoF80 NoF85 NoF90 RITM HRNet32 17 55 151 37 133 336 26 92 336 45 72 135 92 170 501 CDNet ResNet34 67 145 349 91 208 409 132 412 700 108 220 441 193 417 1087 FocalClick HRNet32 78 151 346 131 259 405 116 300 682 111 198 358 289 544 1131 GPCIS ResNet50 27 92 253 84 212 397 34 145 428 31 76 183 141 348 857 SAM-HQ* ViT-B 210 365 582 184 367 539 681 890 1147 495 638 775 642 1179 2055 SimpleClick ViT-B 18 50 137 45 130 317 17 120 345 31 54 93 82 144 405 PiClick** ViT-B 24 65 165 46 131 335 46 194 464 47 95 194 82 160 451 AdaptiveClick ViT-B 24 74 175 81 175 347 65 198 424 27 48 97 57 151 475 MST ViT-B 24 69 209 68 182 370 21 121 362 33 63 146 98 263 789 MFP ViT-B 19 43 144 36 116 314 51 130 378 38 66 158 147 158 491 Ours ViT-B 14 30 113 21 100 305 3 72 299 27 50 100 85 160 422
@@ IV-A Datasets
In this experiment, iSAID and ISPRS Potsdam are adopted as the training datasets, encompassing the majority of instance objects in high-resolution images. During the training phase, images from iSAID and ISPRS Potsdam were cropped to a size of 800×800 pixels, with patches lacking annotation masks excluded. Additionally, the ISPRS Potsdam uses RGB bands. The test datasets employed in the experiments consist of iSAID Val, SandBar, LoveDA Urban, NWPU and WHUBuilding, which differ from the training datasets in terms of imaging conditions, resolution and spectral characteristics. The total number of instances in the above five test datasets is 844, 663, 2049, 851 and 2559, respectively.
TABLE III: Comparison of our method against other fine-tuning approaches in NoC85 and NoF90 on the iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding datasets. The best results are shown in bold. * indicates that the rank of LoRA fine-tuning is 16
iSAID SandBar NWPU LoveDA Urban WHUBuilding NoC85 NoC90 NoC85 NoC90 NoC85 NoC90 NoC85 NoC90 NoC85 NoC90 Pretrained Weights 5.40 8.51 8.69 13.24 4.91 7.39 7.66 10.46 6.38 9.97 Full Fine-Tuning 3.46 6.62 7.91 12.97 4.56 7.30 8.79 11.62 5.63 9.28 LoRA Fine-Tuning * 3.52 6.64 8.14 13.33 4.73 7.41 8.42 11.18 5.23 8.74 Ours 3.38 6.51 7.67 12.90 3.83 6.56 7.62 10.19 5.03 8.42
@@ IV-B Implementation Details
Figure 7: mIoU comparison of our method and other approaches on the iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding datasets. The horizontal axis denotes the number of clicks, while the vertical axis represents the mean IoU.
@@
3) Evaluation Metrics. Adhering to the evaluation strategies proposed by previous works, we adopt average number of clicks (NoC) required to achieve a certain Intersection-over-Union (IoU) ratio. The maximum number of annotated clicks is set to 20. To evaluate performance in high-precision segmentation scenarios, we set the target IoU ratio to 90% and report NoC@80, NoC@85 and NoC@90. Second, we employ NoF, a metric that quantifies the frequency at which an algorithm fails to meet pre-specified performance thresholds with fixed clicks. In the experiments, we set thresholds of 80%, 85% and 90% for NoF. Additionally, we plot the mean Intersection-over-Union (mIoU) scores against the number of clicks.
Figure 8: Segmentation results of RITM(HRNet32), GPCIS(ResNet50), SimpleClick(ViT-B), PiClick(ViT-B), AdaptiveClick(ViT-B), MST(ViT-B), MFP(ViT-B) and our method(ViT-B). Red clicks indicate negative clicks, and green clicks indicate positive clicks. Red regions represent mismatches between the predicted mask and the ground truth mask, while green regions represent matched areas. The red bounding box denotes the ground truth mask boundary.
@@ IV-C Comparative Experiment
1) Comparison of NoC & NoF. As shown in Table I, our RS-ISRefiner methods exhibits strong competitiveness and stability overall across multiple remote sensing image segmentation datasets such as iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding, as well as under different intersection over union (IoU) thresholds including NoC80, NoC85 and NoC90. At lower NoC80 threshold, our method achieves optimal performance on the SandBar, NWPU and LoveDA Urban datasets, and suboptimal performance on the iSAID dataset. However, at the higher IoU thresholds of NoC85 and NoC90, our method attains the best results on all five datasets. Specifically, the average number of clicks required to reach the IoU 85% and IoU 90% thresholds is reduced by 0.352 and 0.242 respectively compared to that of the suboptimal method. This highlights the distinct advantage of our method in refined interactive segmentation, making it suitable for high-quality segmentation and annotation of instance objects in remote sensing images.
Table II quantifies the number of instances that fail to meet the specified segmentation IoU threshold when constrained to 20 clicks across different methods. It is evident that, across the majority of datasets and all IoU thresholds, the count of such unqualified instances using our method is significantly lower than that of comparative approaches. On the NWPU datasets, our method fails to achieve the 80% IoU threshold for merely 3 instances, in contrast to 17 instances for SimpleClick[5] and 51 instances for MFP[18]. Regarding the SandBar dataset with 663 instances, where segmentation poses inheret challenges due to the irregular morphologies and diverse scales of sandbars, our method still manages to satisfy the 90% IoU threshold for 358 instances, outperforming RITM[2](327 instances) and AdaptiveClick[17](316 instances). These results indicate that our method holds tangible advantages in segmentation efficiency under the constraint of limited click counts, which is highly relevant to practical remote sensing image segmentation scenarios.
It is noteworthy that SAM-HQ[20], which adopts a similar technical approach to our method, exhibits the poorest performance among all models in terms of both NoC and NoF metrics. We attribute this to the fact that the backbone of SAM-HQ[20] is pretrained on natural images. Due to the differences between remote sensing images and natural images in aspects such as radiometric energy and spectral accuracy, its strong generalization capability fails to be effectively manifested in remote sensing scenarios. Furthermore, the simplistic design of the adapter module in SAM-HQ[20] renders it incapable of adequately learning and extracting relevant features from remote sensing images, thereby resulting in suboptimal performance. In contrast, the adapter module of our method integrates convolutional networks with the global attention mechanism of Transformer, which can compensate for the deficiency of the backbone in extracting features specific to remote sensing images.
TABLE IV: Efficiency Comparison of Different IIS Methods. denotes testing time per interaction
@@
Ours ViT-B 40.11M 212.49G 48.2ms
TABLE V: Efficiency Comparison of Different Fine-Tuning Methods . denotes testing time per interaction
@@
Ours 40.11M 212.49G 48.2ms
2) Comparison of mIoU Variation. As illustrated in Fig.7, our method performs outstandingly when taking the number of clicks as the independent variable and mIoU as the evaluation metric. Although our method does not achieve high mIoU at the first click, with the increase in the number of clicks, the mIoU curve of our method shows a significant upward trend and approaches a relatively high level. After the final 20th click, the mIoU of our method ranks the best among all methods. Except for the SandBar dataset, the final mIoU of the final mIoU of the iSAID, NWPU, LoveDA Urban and WHUBuilding datasets all exceed 90%; in particular, the iSAID and NWPU datasets can reach an mIoU of over 90% within 10 clicks. Based on the results across the four datasets, our method demonstrates good adaptability and performance advantages, fully verifying the effectiveness and accuracy of the proposed method in click-based segmentation tasks.
@@
As presented in Table III, our proposed adapter fine-tuning method achieves the optimal performance across the NoC85 and NoC90 metrics on five remote sensing datasets: iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding. Compared with the pretrained weights, it reduces the NoC85 score from 5.40 to 3.38 on the iSAID dataset, significantly enhancing the model’s adaptability to the remote sensing domain. Against full fine-tuning, it decreases the NoC85 from 7.91 to 7.67 and NoC90 from 12.97 to 12.90 on the SandBar dataset, maintaining higher segmentation accuracy while reducing parameter updates and avoiding the potential overwrite of pre-trained general visual knowledge by full fine-tuning. Compared with LoRA fine-tuning, it lowers the NoC85 from 4.73 to 3.83 and NoC90 from 7.41 to 6.56 on the NWPU dataset, overcoming the limitation of LoRA’s low-rank structure in expressing remote sensing-specific features (e.g., irregular boundaries, spectral heterogeneity). Ultimately, it balances the preservation of general representations, efficient domain adaptation, and reduced interaction cost, providing an efficient and high-performance fine-tuning solution for interactive remote sensing image segmentation.

### v2

@@ III-C High-Quality Feature Extractor
While the frozen VFM backbone provides rich global semantics and structural priors, these features alone are insufficient to capture fine-grained local variations in remote sensing targets. Therefore, we design a High-Quality Feature Extractor to complement the VFM outputs with detailed spatial representations. This module consists of two components, namely the Image Feature Extractor and the Image Feature Fusion.
Figure 4: High Quality Feature Extractor. This module comprises an image feature extractor (light-colored region in the middle) and a feature fusion module (green region on the right). The image feature extractor includes one ResBlock, three DeformableResBlocks and one 1×1 convolution, with the detailed architectures of ResBlock and Deformable ResBlock shown on the left of the diagram. The feature fusion module is primarily composed of cross-attention blocks, where $\theta$ serves as a learnable parameter.
As shown in Fig.4, the Image Feature Extractor adopts a residual network architecture, comprising one convolutional residual module, three deformable convolutional residual modules and one linear projection layer. Herein, deformable convolutions are introduced to learn sampling position offsets, enabling adaptive adaptation to variations in the shape, orientation and scale of remote sensing targets. Moreover, the integration of the Squeeze-and-Excitation (SE) Attention mechanism into the deformable convolutional residual modules serves as a widely adopted strategy in convolutional networks to enhance the feature extraction capability of the network. Upon processing image $I$ through the Image Feature Extractor, a spatially informative feature $F_{IFE}$ is derived.
@@
Finally, the fused feature $F_{IFE}^{\prime}$ undergoes shape adjustment and is upsampled to a high resolution, yielding the high-quality image feature $F_{HQ}$.
@@ III-D Transformer Decoder
Built upon the high-level semantic features extracted by the frozen Vision Foundation Model, the decoder establishes effective information interaction between the VFM’s global representations and the locally enhanced features from the extractor module. Through bidirectional cross-attention processing, the model establishes efficient information interaction between global semantics and local details, thereby generating prediction results with both global consistency and high-precision boundaries.
As illustrated in Fig.5, First, $F_{mask}$ and $F_{ViT}^{N}$ are concatenated, then undergo channel-wise fusion via the Dense Fusion Convolution, which primarily consists of $1\times 1$ convolution operations, to align dimensions with those of token $Q$. The resulting features $F_{DFC}$, in conjunction with $Q$, are fed into a two-layer bidirectional cross-attention module. Diverging from the conventional Transformer Decoder architecture, we herein reverse the sequence of token self-attention and cross-attention operations. As highlighted in [53], tokens are initially randomly initialized and devoid of image feature information, rendering self-attention operations ineffective for acquiring sufficient information at the initial stage. Thus, the token-to-feature cross-attention operation is performed. $Q$ serves as the Query, with Keys and Values derived from $F_{DFC}$, to align $Q$ with image features. The output then undergoes residual connection and normalization processing. The formula is as follows.
@@ III-E Probability Map Modulation
As noted in [22], probability map modulation enhances or suppresses the probability values of local regions based on the probability map obtained from the previous round and the current user’s click information, making them closer to the foreground or background distribution desired by the user. However, in [22], the modulated local region is a circular area, where both foreground and background pixels may coexist. During the probability modification process, a positive click is intended to enhance the foreground probability, but since there is no distinction between foreground and background within this circular region, the background probability is also inadvertently enhanced. To address this issue, we propose an improved probability map modulation algorithm, as illustrated in Fig.6.
Figure 6: Modulate Map Mechanism. The left image shows the uncorrected result, while the right image presents the corrected version. The modulation window (gray region in the middle image) is extracted using the formula on the left, and the pixel values within this window are refined according to the formula on the right.
After the model outputs the predicted probability map, we obtain the latest click point $u=(x,y,t)$ in the current iteration, where $x$ and $y$ represent the horizontal and vertical coordinates, respectively, and $t$ denotes the click type, with $t=1$ indicating a foreground click and $t=0$ indicating a background click. The modulation radius $R$ is then determined based on the click position: if an opposite-type click exists in the vicinity, $R$ is set to half of the distance from the current click to the nearest opposite-type click; otherwise, a predefined radius $R_{max}$ is used. This design prevents the modulation regions of foreground and background from overlapping. To avoid an excessively small modulation radius, we also introduce a minimum radius constraint $R_{min}$. The equation is given as follows.
@@
where $\Omega$ denotes the set of clicks of opposite type, $\left\|\cdot\right\|$ denotes the Euclidean norm. After obtaining the modulation radius, a circular modulation window is constructed with the click point as the center. Considering that the foreground and background pixel probabilities in the probability map exhibit inherent differences, our method introduces conditional filtering to sieve out irrelevant pixels, extracting only foreground or background regions as the modulation window, which differs from the approach in MFP[22] that modulates the entire circular area. The conditional filtering formula is (9).
@@
In this formula, $\delta$ represents a specific pixel point in the image, $p_{\delta}$ $(\delta\in W)$ represents the probability corresponding to pixel point $\delta$, where $\delta$ is located within the circular modulation region $W$, $mean(\cdot)$ represents the average value of the probability pixels in the region, $mid(\cdot)$ refers to the median of the probability pixels in the region, and $p_{u}$ denotes the pixel probability at the click point. Depending on the click type $t_{u}$, pixels whose probability values fall between the minimum and maximum screening thresholds are retained. Subsequently, the $\gamma$ modulation follows the similar approach of MFP[22]: for different click types, the pixel probabilities in the new modulation region are adjusted, with the adjustment magnitude decreasing as the distance from the click point increases. The relevant formula is (10).
@@
In the formula, $d_{\delta}=\left\|\delta-u\right\|$ represents the distance between the pixel $\delta$ and the click $u$. The modulated probability map will serve as the input for the next iteration of the model. By leveraging the improved probability map modulation algorithm, segmentation errors in target boundary regions can be corrected, thereby enhancing the accuracy and efficiency of interactive segmentation.
@@ IV-A Datasets
In this experiment, iSAID and ISPRS Potsdam are adopted as the training datasets, encompassing the majority of instance objects in high-resolution images. During the training phase, images from iSAID and ISPRS Potsdam were cropped to a size of 800×800 pixels, with patches lacking annotation masks excluded. Additionally, the ISPRS Potsdam uses RGB bands. The test datasets employed in the experiments consist of iSAID Val, SandBar, LoveDA Urban, NWPU and WHUBuilding, which differ from the training datasets in terms of imaging conditions, resolution and spectral characteristics.
TABLE I: Comparison of our method with state-of-the-art IIS approaches in NoC80, NoC85 and NoC90 on the iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding datasets. The best results are highlighted in bold, while the second-best are underlined. * denotes that the backbone parameters of SAM in SAM-HQ are frozen. ** indicates that PiClick was trained without multi-output prediction, using the same model and loss function as the original PiClick.
Method Backbone iSAID SandBar NWPU LoveDA Urban WHUBuilding NoC80 NoC85 NoC90 NoC80 NoC85 NoC90 NoC80 NoC85 NoC90 NoC80 NoC85 NoC90 NoC80 NoC85 NoC90 RITM HRNet32 2.37 3.74 6.71 5.11 8.29 13.54 3.25 4.51 7.17 6.73 8.36 10.82 3.82 5.41 9.00 CDNet ResNet34 5.47 8.14 12.09 8.45 11.76 16.06 5.30 7.84 10.57 9.73 12.58 16.17 7.36 10.20 14.54 FocalClick HRNet32 6.14 8.28 12.14 8.78 11.85 15.29 5.74 7.78 10.70 9.24 11.84 15.14 7.51 10.06 14.24 GPCIS ResNet50 3.78 5.84 9.66 8.35 11.34 15.09 3.87 5.67 8.77 7.00 9.18 12.51 5.48 7.65 11.77 SAM-HQ* ViT-B 10.70 13.47 16.71 11.14 14.98 18.29 11.63 13.31 15.31 15.95 17.71 19.15 11.23 14.40 18.13 SimpleClick ViT-B 2.40 3.76 6.66 5.25 8.22 13.34 3.08 4.57 7.12 6.82 8.33 10.62 3.64 5.08 8.42 PiClick** ViT-B 3.01 4.61 8.28 5.42 8.64 14.06 4.12 6.06 8.61 8.31 10.16 12.85 3.87 5.35 8.98 AdaptiveClick ViT-B 2.89 4.61 7.79 6.29 9.37 13.76 3.72 5.36 7.81 6.84 8.44 10.67 3.54 5.03 8.46 MST ViT-B 2.77 4.40 8.20 6.15 9.55 14.19 3.08 4.75 7.71 6.84 8.72 11.60 4.16 6.17 10.55 MFP ViT-B 2.20 3.46 6.62 4.63 7.91 12.97 3.24 4.56 7.30 7.13 8.79 11.62 4.01 5.63 9.28 $\text{VFM-ISRefiner}_{\text{(ours)}}$ ViT-B 2.25 3.38 6.51 4.54 7.67 12.90 2.53 3.83 6.56 5.99 7.62 10.19 3.65 5.03 8.42
TABLE II: Comparison of our method against other fine-tuning approaches in NoC85 and NoC90 on the iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding datasets. The best results are shown in bold. * indicates that the rank of LoRA fine-tuning is 16
iSAID SandBar NWPU LoveDA Urban WHUBuilding NoC85 NoC90 NoC85 NoC90 NoC85 NoC90 NoC85 NoC90 NoC85 NoC90 Pretrained Weights 5.40 8.51 8.69 13.24 4.91 7.39 7.66 10.46 6.38 9.97 Full Fine-Tuning 3.46 6.62 7.91 12.97 4.56 7.30 8.79 11.62 5.63 9.28 LoRA Fine-Tuning * 3.52 6.64 8.14 13.33 4.73 7.41 8.42 11.18 5.23 8.74 $\text{VFM-ISRefiner}_{\text{(ours)}}$ 3.38 6.51 7.67 12.90 3.83 6.56 7.62 10.19 5.03 8.42
@@ IV-B Implementation Details
Figure 7: mIoU comparison of our method and other approaches on the iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding datasets. The horizontal axis denotes the number of clicks, while the vertical axis represents the mean IoU. In the legend, ”RITM-HRNET32[89.94]” denotes the RITM model with HRNET32 backbone, which achieves a maximum mIoU value of 89.94% on the dataset.
@@
3) Evaluation Metrics. Adhering to the evaluation strategies proposed by previous works, we adopt average number of clicks (NoC) required to achieve a certain Intersection-over-Union (IoU) ratio. The maximum number of annotated clicks is set to 20. To evaluate performance in high-precision segmentation scenarios, we set the target IoU ratio to 90% and report NoC80, NoC85 and NoC90. Additionally, we plot the mean Intersection-over-Union (mIoU) scores against the number of clicks.
Figure 8: Segmentation results of RITM(HRNet32), GPCIS(ResNet50), SimpleClick(ViT-B), PiClick(ViT-B), AdaptiveClick(ViT-B), MST(ViT-B), MFP(ViT-B) and our VFM-ISRefiner(ViT-B). Red clicks indicate negative clicks, and green clicks indicate positive clicks. Red regions represent mismatches between the predicted mask and the ground truth mask, while green regions represent matched areas. The red bounding box denotes the ground truth mask boundary.
@@ IV-C Comparative Experiment
1) Comparison of NoC. As shown in Table I, our VFM-ISRefiner method exhibits strong competitiveness and stability overall across multiple remote sensing image segmentation datasets such as iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding, as well as under different intersection over union (IoU) thresholds including NoC80, NoC85 and NoC90. At lower NoC80 threshold, our method achieves optimal performance on the SandBar, NWPU and LoveDA Urban datasets, and suboptimal performance on the iSAID dataset. However, at the higher IoU thresholds of NoC85 and NoC90, our method attains the best results on all five datasets. Specifically, the average number of clicks required to reach the IoU 85% and IoU 90% thresholds is reduced by 0.352 and 0.242 respectively compared to that of the suboptimal method. This highlights the distinct advantage of our method in refined interactive segmentation, making it suitable for high-quality segmentation and annotation of instance objects in remote sensing images.
It is noteworthy that SAM-HQ[24], which adopts a similar technical approach to our method, exhibits the poorest performance among all models in terms of NoC metric. We attribute this to the fact that the backbone of SAM-HQ[24] is pretrained on natural images. Due to the differences between remote sensing images and natural images in aspects such as radiometric energy and spectral accuracy, its strong generalization capability fails to be effectively manifested in remote sensing scenarios. Furthermore, the simplistic design of the adapter module in SAM-HQ[24] renders it incapable of adequately learning and extracting relevant features from remote sensing images, thereby resulting in suboptimal performance. In contrast, the adapter module of our method integrates convolutional networks with the global attention mechanism of Transformer, which can compensate for the deficiency of the backbone in extracting features specific to remote sensing images. In addition, the improved probability map modulation method proposed in this paper can correct the segmentation errors of target boundaries, thereby further improving the segmentation accuracy of our model.
TABLE III: Efficiency Comparison of Different IIS Methods. denotes testing time per interaction
@@
$\text{VFM-ISRefiner}_{\text{(ours)}}$ ViT-B 40.11M 212.49G 48.2ms
TABLE IV: Efficiency Comparison of Different Fine-Tuning Methods . denotes testing time per interaction
@@
$\text{VFM-ISRefiner}_{\text{(ours)}}$ 40.11M 212.49G 48.2ms
2) Comparison of mIoU Variation. As illustrated in Fig.7, we adopt the number of clicks as the independent variable and mIoU as the evaluation metric to conduct a comparative performance analysis of competing methods; specifically, while the initial mIoU of our proposed method at the first click is not competitive, its mIoU curve exhibits a steep upward trend with increasing click counts, rising sharply within the first 5-10 clicks as reflected in the subfigures corresponding to each dataset, and then rapidly approaches a high-performance plateau that stabilizes after 10-15 clicks. After the final 20th click, our method attains the optimal mIoU among all competing methods on four datasets, namely iSAID (92.67%), SandBar (88.93%), NWPU (93.87%), and WHUBuilding (91.59%), and it still achieves the second-highest mIoU with a value of 93.3% on the LoveDA Urban dataset. Except for the SandBar dataset, the final mIoU values of iSAID, NWPU, LoveDA Urban, and WHUBuilding all surpass 90%, and notably, the iSAID and NWPU datasets, where our method reaches the 90% mIoU threshold within 8 and 10 clicks, respectively demonstrate that our method is capable of rapidly converging to a high mIoU level with a small number of interaction clicks, a characteristic clearly reflected in the curve trends of these two datasets. Collectively, the experimental results across the five datasets indicate that our method possesses robust adaptability and prominent performance advantages in terms of both rapid convergence and high final accuracy, thereby fully validating the effectiveness and precision of the proposed method in click-based interactive segmentation tasks.
@@
3) Comparison With Fine-Tune Methods. As presented in Table II, the first row presents the experimental results using the MFP weights pretrained on natural images. The second row shows the results of MFP with full-parameter fine-tuning on the remote sensing dataset proposed in this paper. The third row denotes the experimental results of LoRA fine-tuning applied to the ViT backbone of MFP. The fourth row reports the results obtained by the fine-tuning method proposed in this study. Our proposed adapter fine-tuning method achieves the optimal performance across the NoC85 and NoC90 metrics on five remote sensing datasets: iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding. Compared with the pretrained weights, it reduces the NoC85 score from 5.40 to 3.38 on the iSAID dataset, significantly enhancing the model’s adaptability to the remote sensing domain. Against full fine-tuning, it decreases the NoC85 from 7.91 to 7.67 and NoC90 from 12.97 to 12.90 on the SandBar dataset, maintaining higher segmentation accuracy while reducing parameter updates and avoiding the potential overwrite of pretrained general visual knowledge by full fine-tuning. Compared with LoRA fine-tuning, it lowers the NoC85 from 4.73 to 3.83 and NoC90 from 7.41 to 6.56 on the NWPU dataset, overcoming the limitation of LoRA’s low-rank structure in expressing remote sensing-specific features (e.g., irregular boundaries, spectral heterogeneity). Ultimately, it balances the preservation of general representations, efficient domain adaptation, and reduced interaction cost, providing an efficient and high-performance fine-tuning solution for interactive remote sensing image segmentation.